OpenStudy (anonymous):

A 4.0kg object traveling westward at 25ms^-1 hits a 15kg object at rest. The 4.0kg object bounces eastward at 8.0ms^-1. What is the speed and directions of the 15kg object?

OpenStudy (anonymous):

Pi=Pf M1V1+M2V2=M1V1`+M2V2` 4*25+0=4*8+15V2` 4.53=V2` westward